OpenStudy (anonymous):

Evaluate the following expression using the values given. Find x + 2y if x = 3 and y = 4

OpenStudy (anonymous):

We need to plug in 3 for the x and 4 for the y: \[x+2y=3+2(4)=3+8=11\]